# Satisficing Regret Minimization in Bandits

### v2

@@ 3 SELECT: Satisficing Regret Minimization via Sampling and Lower Confidence Bound Testing
Step 2. One challenge from using the LCB test in Step 3 is that it is a more conservative design compared to prior works. In fact, if we directly enter Step 3 without Step 2, the LCB of $r(\hat{X_{i}})$ can easily fall below $S$ even if $r(\hat{X}_{i})\geq S$. This is because we may not have pulled $\hat{X}_{i}$ for enough times in Step 1 and the corresponding confidence interval might be large. As a result, the lower confidence bound of $r(\hat{X}_{i})$, which is the difference between empirical mean and confidence radius, can be significantly below $S$. This indicates, without the forced sampling in Step 2, major satisficing regret can be incurred due to frequent re-start of a new round. With the forced sampled data collected from Step 2, SELECT ensures that the width of the confidence interval is of order $\tilde{O}(T_{i}^{-1/2})=\tilde{O}(\gamma_{i})$ (i.e., same as $\mathbb{E}[r(X^{*})-r(\hat{X}_{i})]$). Consequently, whenever $\gamma_{i}$ shrinks to well below $\Delta_{S}^{*}$, SELECT gradually becomes less likely to terminate a round (see the forthcoming Proposition 2);
Step 3. In Step 3, SELECT compares the LCB of $\hat{X_{i}}$’s mean reward against $S$ to determine if $X_{i}$ is a satisficing arm. This deviates from prior works that use UCB (Garivier et al. 2019) or empirical mean (Michel et al. 2023). However, it turns out that this is essential in achieving a constant satisficing regret that does not scale with $1/\Delta_{S}$, which can quickly explode in many cases. As we will see, once entering Step 3, SELECT can terminate a round within 1 additional time step in expectation when facing a non-satisficing arm (i.e., (3) in the proof of the forthcoming Proposition 1). In contrast, if it follows prior works to use UCB or empirical mean instead, the number of time steps required (and hence, the satisficing regret) will unavoidably scale with $1/\Delta_{S}$ (see e.g., Theorem 9 of Garivier et al. 2019 or Theorem 1 of Michel et al. 2023).
@@
$\gamma_{i}\leftarrow 2^{-i(1-\alpha)/\alpha}$ for all $i=1,2,3\dots$
@@
     Run $\texttt{ALG}(t_{i})$, let $\{\bar{X}_{s}\}_{s\in[t_{i}]}$ be the trajectory of arm pulls
     Sample $R$ uniformly at random from $\{1,2,\dots,t_{i}\}$ and set $\hat{X}_{i}\leftarrow\bar{X}_{R}$
@@
     Let $t$ be the current time step, pull arm $\hat{X}_{i}$ for the next $T_{i}$ time steps, and set $\hat{r}_{i}^{\text{tot}}\leftarrow\sum_{s=t}^{t+T_{i}-1}Y_{s}$
     while $LCB(\hat{X}_{i})\geq S$ do
@@
         Pull arm $\hat{X}_{i}$ and observe $Y_{t}$, where $t$ is the current time step
@@
$LCB(\hat{X}_{i})\leftarrow\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}$
@@ 4 Regret Analysis
In this section, we analyze the regret bounds of SELECT. We begin by providing an upper bound of satisficing regret in the realizable case, i.e., $r(X^{*})\geq S$.
@@
If $r(X^{*})\geq S$, then the satisficing regret of SELECT is bounded by
$\texttt{Regret}_{S}\leq\min\left\{C_{1}^{\frac{1}{1-\alpha}}\left(\dfrac{1}{\Delta_{S}^{*}}\right)^{\frac{\alpha}{1-\alpha}}\cdot\textup{polylog}\left(\dfrac{C_{1}}{\Delta_{S}^{*}}\right),\,C_{1}T^{\alpha}\cdot\textup{polylog}(T)\right\}.$ (1)
We remark that, from the first term on the RHS of (1), SELECT can achieve a constant (w.r.t. $T$) satisficing regret in the realizable case. Moreover, even when $T$ is relatively small and the constant satisficing regret guarantee is loose compared to the oracle’s regret bound, it is able to adapt to the oracle’s performance.
@@
A complete proof for Theorem 1 is provided in Section A. The proof relies on two critical results. In the first one, we show an upper bound on the satisficing regret of round $i$.
@@
If $r(X^{*})\geq S$, then the satisficing regret incurred in round $i$ is bounded by
@@
If $r(X^{*})\geq S$, then for every $i$ that satisfies
@@
The proof of this proposition is provided in Section A.2 of the appendix. Since each round of SELECT runs independently, Proposition 2 indicates that the total number of rounds for SELECT can be upper-bounded by a shifted geometric random variable with success rate $3/4$. Combining the results of Proposition 1 and Proposition 2 we are able to establish the statement in Theorem 1. ∎
@@
If $r(X^{*})<S$, the standard regret of SELECT is bounded by
$\texttt{Regret}\leq C_{1}T^{\alpha}\cdot\textup{polylog}(T).$
With the above results, we establish that under a very general condition, SELECT does achieve constant satisficing regret in the realizable case without compromising the standard regret guarantee in the non-realizable case. Altogether, these empower SELECT the potential to become a general algorithm for decision-making under uncertainty.
@@ 5 Examples
In what follows, we instantiate SELECT to several popular bandit optimization settings, including finite-armed bandits, concave bandits, and Lipschitz bandits. Along the way, we showcase how our results enable constant satisficing regret for bandits with large and even infinite arm space, which was not achieved by existing algorithms. We remark that the examples given here are non-exhaustive and similar results can be derived for other bandit optimization settings such as linear bandits. In all three examples of this section, we assume the mean reward of each arm is in $[0,1]$.
1. Finite-Armed Bandits: Consider a finite-armed bandit problem with $K$ arms, i.e., $\mathcal{X}=[K]$. In this case, both the UCB algorithm (see, e.g., Bubeck and Cesa-Bianchi 2012) and Thompson sampling (see, e.g., Agrawal and Goyal 2017) achieve a regret bound of $O(\sqrt{KT\log(T)})$. Combining them with Theorems 1 and 2, we have the following corollary. At the end of this section, we also provide some discussions on the lower bounds of the satisficing regrets.
@@
By using either the UCB algorithm or Thompson sampling as ALG, if $r(X^{*})\geq S$, the satisficing regret of SELECT is bounded by
$\texttt{Regret}_{S}\leq\min\left\{\dfrac{K}{\Delta_{S}^{*}}\cdot\textup{polylog}\left(\dfrac{K}{\Delta_{S}^{*}}\right),\,\sqrt{KT}\cdot\textup{polylog}(T)\right\};$
If $r(X^{*})<S$, the standard regret of SELECT is bounded by $\texttt{Regret}\leq\sqrt{KT}\cdot\textup{polylog}(T).$
@@
Garivier et al. (2019), Michel et al. (2023) also provide algorithms for satisficing regret for finite-armed bandits. While our satisficing regret bound is incomparable to Garivier et al. (2019), which attains a $O(K/\Delta_{S})$ satisficing regret, we provide major improvement over Michel et al. (2023), which achieves $O(K/\Delta_{S}+K/(\Delta_{S}^{*})^{2})$ satisficing regret. Compared to Michel et al. (2023), our regret bound remove the dependence on $\Delta_{S}$ and the additional $1/\Delta_{S}^{*}$ factor. We also point out that if we want to achieve the same satisficing regret as Garivier et al. (2019), we can simply change the LCB test in Step 3 to a UCB test.
2. Concave Bandits: Consider a concave bandit setting, i.e., $\mathcal{X}\subset\mathbb{R}^{d}$ is a bounded convex set, and $r(X)$ is a concave and $1$-Lipschitz continuous function defined on $\mathcal{X}$. Agarwal et al. (2011) gives an algorithm that enjoys $\text{poly}(d)\sqrt{T}\cdot\text{polylog}(T)$ regret. Together with Theorems 1 and 2, we have the following corollary.
@@
By using the algorithm in Agarwal et al. (2011) as ALG, if $r(X^{*})\geq S$, then the satisficing regret of SELECT is bounded by
$\texttt{Regret}_{S}\leq\min\left\{\dfrac{\textup{poly}(d)}{\Delta_{S}^{*}}\cdot\textup{polylog}\left(\dfrac{d}{\Delta_{S}^{*}}\right),\textup{poly}(d)\sqrt{T}\cdot\textup{polylog}(T)\right\};$
If $r(X^{*})<S$, the standard regret of SELECT is bounded by $\texttt{Regret}\leq\textup{poly}(d)\sqrt{T}\cdot\textup{polylog}(T).$
3. Lipschitz Bandits: Consider a Lipschitz bandit setting in $d$ dimensions, i.e., $\mathcal{X}=[0,1]^{d}$, and $r(X)$ is an $L$-Lipschitz function (here Lipschitz functions are defined in the sense of $\infty$-norm). Bubeck et al. (2011) introduces a uniformly discretized UCB algorithm to achieve a $O(L^{d/(d+2)}T^{(d+1)/(d+2)}\sqrt{\log(T)})$ regret upper bound bound. Together with Theorem 1 and Theorem 2, we have the following corollary.
@@
By using the UCB algorithm with uniform discretization in Bubeck et al. (2011) as ALG, if $r(X^{*})\geq S$, then the satisficing regret of SELECT is bounded by
$\texttt{Regret}_{S}\leq\min\left\{\dfrac{L^{d}}{(\Delta_{S}^{*}/2)^{d+1}}\cdot\textup{polylog}\left(\dfrac{L}{\Delta_{S}^{*}}\right),\,L^{\frac{d}{d+2}}T^{\frac{d+1}{d+2}}\cdot\textup{polylog}(T)\right\};$
If $r(X^{*})<S$, the standard regret of SELECT is bounded by $\texttt{Regret}\leq L^{\frac{d}{d+2}}T^{\frac{d+1}{d+2}}\cdot\textup{polylog}(T).$
@@
Recall that $\Delta_{S}=\min\{S-r(X):r(X)<S\}$, one can easily verify that in the realizable case, $\Delta_{S}=0$ for both concave bandits and the Lipschitz bandits. As such, one cannot directly apply the results in Garivier et al. (2019), Michel et al. (2023) to acquire a constant satisficing regret. With the notion of exceeding gap $\Delta_{S}^{*}$ and SELECT, we establish constant satisficing regret bounds for these two settings.
4. Lower Bounds: To complement our main results, we also present the satisficing regret lower bounds for finite-armed bandits and concave bandits. The first one is a satisficing regret lower bound for the finite-armed bandits.
@@
For every non-anticipatory learning algorithm $\pi$, every $\Delta>0$ and $T\geq 1/\Delta^{2}$, there exists an instance of two-armed bandit such that $r(X^{*})-S=\Delta$, and the satisficing regret incurred by $\pi$ is at least $\Omega(1/\Delta)$.
@@
In Michel et al. (2023), the authors adapt the results from Bubeck et al. (2013) (for MAB with known optimal mean reward) to establish a $\Omega(1/\Delta_{S})$ satisficing regret lower bound for finite-armed bandits. This is different than the one in Theorem3, which is $\Omega(1/\Delta)$ with $\Delta$ being the exceeding gap (i.e., $\Delta_{S}^{*}$). This difference originates from the lower bound instances. The lower bound instance in Bubeck et al. (2013) (when adapted to satisficing bandits) sets the threshold value to the optimal mean reward, i.e., $S=r(X^{*})\,,$ which makes both the satisficing regret and standard regret lower bounds to be $\Omega(1/\Delta_{S})$. In our lower bound proof, we allow $S$ to be smaller than $r(X^{*})$, which leads to a lower bound that depends on the difference between these two quantities.
Next, we provide a satisficing regret lower bound for bandits with concave rewards.

### v3

@@ 3 SELECT: Satisficing Regret Minimization via Sampling and Lower Confidence Bound Testing
Step 2. One challenge from using the LCB test in Step 3 is that it is a more conservative design compared to prior works. In fact, if we directly enter Step 3 without Step 2, the LCB of $r(\hat{A_{i}})$ can easily fall below $S$ even if $r(\hat{A}_{i})\geq S$. This is because we may not have pulled $\hat{A}_{i}$ for sufficiently many times in Step 1 and the corresponding confidence interval might be large. As a result, the lower confidence bound of $r(\hat{A}_{i})$, which is the difference between empirical mean and confidence radius, can be significantly below $S$. This indicates, without the forced sampling in Step 2, major satisficing regret can be incurred due to frequent re-start of a new round. With the forced sampled data collected from Step 2, SELECT ensures that the width of the confidence interval is of order $\tilde{O}(T_{i}^{-1/2})=\tilde{O}(\gamma_{i})$ (i.e., same as $\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]$). Consequently, whenever $\gamma_{i}$ shrinks to well below $\Delta_{S}^{*}$, SELECT gradually becomes less likely to terminate a round (see the forthcoming Proposition 2);
Step 3. In Step 3, SELECT compares the LCB of $\hat{A_{i}}$’s mean reward against $S$ to determine if $A_{i}$ is a satisficing arm. This deviates from prior works that use UCB (Garivier et al. 2019) or empirical mean (Michel et al. 2023). However, it turns out that this is essential in achieving a constant expected satisficing regret that does not scale with $1/\Delta_{S}$, which can quickly explode in many cases. As we will see, once entering Step 3, SELECT can terminate a round within 1 additional time step in expectation when facing a non-satisficing arm (i.e., (6) in the proof of the forthcoming Proposition 1). In contrast, if it follows prior works to use UCB or empirical mean instead, the number of time steps required (and hence, the satisficing regret) will unavoidably scale with $1/\Delta_{S}$ (see e.g., Theorem 9 of Garivier et al. 2019 or Theorem 1 of Michel et al. 2023).
@@
Set $\gamma_{i}\leftarrow 2^{-i(1-\alpha)/\alpha}$ for all $i=1,2,3\dots$
@@
     Run $\texttt{ALG}(t_{i})$, let $\{\bar{A}_{s}\}_{s\in[t_{i}]}$ be the trajectory of arm pulls
     Sample $q$ uniformly at random from $\{1,2,\dots,t_{i}\}$ and set $\hat{A}_{i}\leftarrow\bar{A}_{q}$
@@
     Let $t$ be the current time step, pull arm $\hat{A}_{i}$ for the next $T_{i}$ time steps, and set $\hat{r}_{i}^{\text{tot}}\leftarrow\sum_{s=t}^{t+T_{i}-1}Y_{s}$
     while $LCB(\hat{A}_{i})\geq S$ do
@@
         Pull arm $\hat{A}_{i}$ and observe $Y_{t}$, where $t$ is the current time step
@@
$LCB(\hat{A}_{i})\leftarrow\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}$
@@ 4 Regret Analysis
In this section, we analyze the regret bounds of SELECT. We begin by providing an upper bound of expected satisficing regret in the realizable case, i.e., $r(A^{*})\geq S$.
@@
If $r(A^{*})\geq S$, then the expected satisficing regret of SELECT is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\min\left\{C_{1}^{\frac{1}{1-\alpha}}\left(\dfrac{1}{\Delta_{S}^{*}}\right)^{\frac{\alpha}{1-\alpha}}\cdot\textup{polylog}\left(\dfrac{C_{1}}{\Delta_{S}^{*}}\right),\,C_{1}T^{\alpha}\cdot\textup{polylog}(T)\right\}.$ (1)
We remark that, from the first term on the RHS of (1), SELECT can achieve a constant (w.r.t. $T$) expected satisficing regret in the realizable case. Moreover, even when $T$ is relatively small and the constant expected satisficing regret guarantee is loose compared to the oracle’s regret bound, it is able to adapt to the oracle’s performance and achieve a sub-linear in $T$ bound on expected satisficing regret.
@@
A complete proof for Theorem 1 is provided in Section A. The proof relies on two critical results. In the first one, we show an upper bound on the expected satisficing regret of round $i$.
@@
If $r(A^{*})\geq S$, then the expected satisficing regret incurred in round $i$ is bounded by
@@
If $r(A^{*})\geq S$, then for every $i$ that satisfies
$\dfrac{32C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\max\{\beta,1/2\}}\leq\Delta_{S}^{*}\,,$ (2)
@@
The proof of this proposition is provided in Section A.2 of the appendix. Since each round of SELECT runs independently, Proposition 2 indicates that the total number of rounds for SELECT can be upper-bounded by a shifted geometric random variable with success rate $3/4$.
With Proposition 1 and Proposition 2, we are ready to finish the proof of Theorem 1. We denote $i_{0}$ as the minimum integer such that (2) holds. We prove that $2^{i_{0}(1-\alpha)/\alpha}=1/\gamma_{i_{0}}=C_{1}/\Delta_{S}^{*}\cdot\text{polylog}(C_{1}/\Delta_{S}^{*})$. We further have $i_{0}=\Theta(\log(C_{1}/\Delta_{S}^{*}))$. Therefore by Proposition 1 the expected satisficing regret incurred in the first $i_{0}$ rounds is bounded by
$\sum_{i=1}^{i_{0}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\frac{\alpha}{1-\alpha}}\log(2/\gamma_{i})^{\beta}\leq C_{1}^{\frac{1}{1-\alpha}}\left(\dfrac{1}{\Delta_{S}^{*}}\right)^{\frac{\alpha}{1-\alpha}}\cdot\text{polylog}\left(\dfrac{C_{1}}{\Delta_{S}^{*}}\right).$
Then we bound the expected satisficing regret incurred after round $i_{0}$. By Proposition 2 SELECT enters round $i_{0}+k$ with probability $4^{-k}$. Therefore the expected satisficing regret after round $i_{0}$ is
$\sum_{k=1}^{\infty}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i_{0}+k}^{-\frac{\alpha}{1-\alpha}}\log(2/\gamma_{i_{0}+k})^{\beta}\cdot 4^{-k}=\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i_{0}}^{-\frac{\alpha}{1-\alpha}}\sum_{k=1}^{\infty}2^{-k}\left(\dfrac{(1-\alpha)(i_{0}+k)}{\alpha}+1\right)^{\beta}\log(2)^{\beta}.$
We have shown that $\gamma_{i_{0}}^{-\alpha/(1-\alpha)}=(C_{1}/\Delta_{S}^{*})^{\alpha/(1-\alpha)}\cdot\text{polylog}(C_{1}/\Delta_{S}^{*})$. We further show that the summation above is polynomial in $i_{0}$, thus polynomial in $\log(C_{1}/\Delta_{S}^{*})$. Therefore the expected satisficing regret is upper bounded by $C_{1}^{1/(1-\alpha)}(1/\Delta_{S}^{*})^{\alpha/(1-\alpha)}\cdot\text{polylog}(C_{1}/\Delta_{S}^{*})$. Therefore the expected satisficing regret is upper bounded by the first term in the regret bound.
We further show that the expected satisficing regret is upper bounded by the second term in the regret bound. Let $i_{\max}$ be the total number of rounds for SELECT. We show that $t_{i_{\max}}=\gamma_{i_{\max}}^{-1/(1-\alpha)}=O(T)$, thus $i_{\max}=O(\log(T))$. By Proposition 1 the expected satisficing regret of each round is upper bounded by $C_{1}T^{\alpha}\cdot\text{polylog}(T)$. Since the total number of rounds is upper bounded by $O(\log(T))$, the expected satisficing regret of SELECT is also upper bounded by $C_{1}T^{\alpha}\cdot\text{polylog}(T)$. ∎
@@
If $r(A^{*})<S$, the standard regret of SELECT is bounded by
$\mathbb{E}[\texttt{Regret}]\leq C_{1}T^{\alpha}\cdot\textup{polylog}(T).$
Proof of Theorem 2 is provided in Appendix B. Theorem 2 implies that the expected standard regret of SELECT is sub-linear in $T$ in the non-realizable case. The expected standard regret bound of SELECT matches that of the learning oracle within logarithmic factors of $T$, indicating a similar performance guarantee of SELECT compared to the standard regret minimization algorithm in the corresponding class of bandit optimization problems.
With the above results, we establish that under a very general condition, SELECT does achieve constant expected satisficing regret in the realizable case without compromising the standard regret guarantee in the non-realizable case. Altogether, these empower SELECT the potential to become a general algorithm for decision-making under uncertainty.
@@ 5 Examples
In what follows, we instantiate SELECT to several popular classes of bandit optimization problems, including finite-armed bandits, concave bandits, and Lipschitz bandits. Along the way, we showcase how our results enable constant expected satisficing regret for bandits with large and even infinite arm set, which was not achieved by existing algorithms. We remark that the examples given here are non-exhaustive and similar results can be derived for other classes of bandit optimization problems such as linear bandits. In all three examples of this section, we assume the mean reward of each arm is in $[0,1]$.
1. Finite-Armed Bandits: Consider any instance of finite-armed bandit with $K$ arms, i.e., $\mathcal{A}=[K]$. In this case, both the UCB algorithm (see, e.g., Bubeck and Cesa-Bianchi 2012) and Thompson sampling (see, e.g., Agrawal and Goyal 2017) achieve a regret bound of $O(\sqrt{KT\log(T)})$. Combining them with Theorems 1 and 2, we have the following corollary. At the end of this section, we also provide some discussions on the lower bounds of the expcted satisficing regret.
@@
By using either the UCB algorithm or Thompson sampling as ALG, if $r(A^{*})\geq S$, the expected satisficing regret of SELECT is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\min\left\{\dfrac{K}{\Delta_{S}^{*}}\cdot\textup{polylog}\left(\dfrac{K}{\Delta_{S}^{*}}\right),\,\sqrt{KT}\cdot\textup{polylog}(T)\right\};$
If $r(A^{*})<S$, the standard regret of SELECT is bounded by $\mathbb{E}[\texttt{Regret}]\leq\sqrt{KT}\cdot\textup{polylog}(T).$
@@
Garivier et al. (2019), Michel et al. (2023) also provide algorithms for the expected satisficing regret in finite-armed bandit settings. While our expected satisficing regret bound is incomparable to Garivier et al. (2019), which attains a $O(K/\Delta_{S})$ expected satisficing regret, we provide major improvement over Michel et al. (2023), which achieves $O(K/\Delta_{S}+K/(\Delta_{S}^{*})^{2})$ expected satisficing regret. Compared to Michel et al. (2023), our regret bound remove the dependence on $\Delta_{S}$ and the additional $1/\Delta_{S}^{*}$ factor. We also point out that if we want to achieve the same expected satisficing regret as Garivier et al. (2019), we can simply change the LCB test in Step 3 to a UCB test.
2. Concave Bandits: Consider any instance of the concave bandit, i.e., $\mathcal{A}\subset\mathbb{R}^{d}$ is a bounded convex set, and $r(A)$ is a concave and $1$-Lipschitz continuous function defined on $\mathcal{A}$. Agarwal et al. (2011) gives an algorithm that enjoys $\text{poly}(d)\sqrt{T}\cdot\text{polylog}(T)$ regret. Together with Theorems 1 and 2, we have the following corollary.
@@
By using the algorithm in Agarwal et al. (2011) as ALG, if $r(A^{*})\geq S$, then the expected satisficing regret of SELECT is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\min\left\{\dfrac{\textup{poly}(d)}{\Delta_{S}^{*}}\cdot\textup{polylog}\left(\dfrac{d}{\Delta_{S}^{*}}\right),\textup{poly}(d)\sqrt{T}\cdot\textup{polylog}(T)\right\};$
If $r(A^{*})<S$, the standard regret of SELECT is bounded by $\mathbb{E}[\texttt{Regret}]\leq\textup{poly}(d)\sqrt{T}\cdot\textup{polylog}(T).$
3. Lipschitz Bandits: Consider any instance of the Lipschitz bandit in $d$ dimensions, i.e., $\mathcal{A}=[0,1]^{d}$, and $r(A)$ is an $L$-Lipschitz function (here Lipschitz functions are defined in the sense of $\infty$-norm). Bubeck et al. (2011) introduces a uniformly discretized UCB algorithm to achieve a $O(L^{d/(d+2)}T^{(d+1)/(d+2)}\sqrt{\log(T)})$ regret upper bound bound. Together with Theorem 1 and Theorem 2, we have the following corollary.
@@
By using the UCB algorithm with uniform discretization in Bubeck et al. (2011) as ALG, if $r(A^{*})\geq S$, then the expected satisficing regret of SELECT is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\min\left\{\dfrac{L^{d}}{(\Delta_{S}^{*}/2)^{d+1}}\cdot\textup{polylog}\left(\dfrac{L}{\Delta_{S}^{*}}\right),\,L^{\frac{d}{d+2}}T^{\frac{d+1}{d+2}}\cdot\textup{polylog}(T)\right\};$
If $r(A^{*})<S$, the standard regret of SELECT is bounded by $\mathbb{E}[\texttt{Regret}]\leq L^{\frac{d}{d+2}}T^{\frac{d+1}{d+2}}\cdot\textup{polylog}(T).$
@@
Recall that $\Delta_{S}=\min\{S-r(A):r(A)<S\}$, one can easily verify that in the realizable case, $\Delta_{S}=0$ for both concave bandits and the Lipschitz bandits. As such, one cannot directly apply the results in Garivier et al. (2019), Michel et al. (2023) to acquire a constant expected satisficing regret. With the notion of exceeding gap $\Delta_{S}^{*}$ and SELECT, we establish constant expected satisficing regret bounds for these two classes of bandit optimization problems.
4. Lower Bounds: To complement our main results, we also present the expected satisficing regret lower bounds for finite-armed bandits and concave bandits. The first one is an expected satisficing regret lower bound for the finite-armed bandits.
@@
For every non-anticipatory learning algorithm $\pi$, every $\Delta>0$ and $T\geq 1/\Delta^{2}$, there exists an instance of two-armed bandit such that $r(A^{*})-S=\Delta$, and the expected satisficing regret incurred by $\pi$ is at least $\Omega(1/\Delta)$.
Theorem 3 establishes a $\Omega(1/\Delta_{S}^{*})$ lower bound on expected satisficing regret for finite-armed bandits. The lower bound in Theorem 3 matches the expected satisficing regret bound of SELECT in Corollary 1 up to a factor of $K$ and a logarithmic factor of $K/\Delta_{S}^{*}$. If we consider the number of arms $K$ as a constant, then the expected satisficing regret bound of SELECT for finite-armed bandits is near-optimal.
@@
In Michel et al. (2023), the authors adapt the results from Bubeck et al. (2013) (for MAB with known optimal mean reward) to establish a $\Omega(1/\Delta_{S})$ expected satisficing regret lower bound for finite-armed bandits. This is different than the one in Theorem3, which is $\Omega(1/\Delta)$ with $\Delta$ being the exceeding gap (i.e., $\Delta_{S}^{*}$). This difference originates from the lower bound instances. The lower bound instance in Bubeck et al. (2013) (when adapted to satisficing bandits) sets the threshold value to the optimal mean reward, i.e., $S=r(A^{*})\,,$ which makes both the expected satisficing regret and standard regret lower bounds to be $\Omega(1/\Delta_{S})$. In our lower bound proof, we allow $S$ to be smaller than $r(A^{*})$, which leads to a lower bound that depends on the difference between these two quantities.
Next, we provide an expected satisficing regret lower bound for bandits with concave rewards.